# Logical perspectives on learning statistical objects

### v1

@@ 6.4 An Alternate Approach to Realizable Online Learning
Thus far we have shown that realizable online learning is not preserved under moving to statistical classes. We have indicated that this is related to another pathology, that online learnability is not preserved under applying continuous mappings. We will now look at using this intuition to “fix” the pathologies of realizable online learning. We will do this by changing the definition, applying alternate loss functions which do not sum the cumulative losses, but rather discretize them. The change is motivated by similar discretizations that appear in continuous logic.
@@
Definition 17 (Online learnability for a general loss function).
@@
Proof.
The same learning algorithm will suffice. On any given run of the algorithm, the regret with $\ell_{1}$ as the loss function will be at most the regret with $\ell_{2}$ as the loss function:
@@
so regret with loss function $\ell_{1}$ will satisfy the same upper bound required of regret with loss function $\ell_{2}$, up to the fixed factor $C$. ∎
@@
Definition 18 ($\epsilon$-truncated loss functions).
@@
Proof.
Online learnability is equivalent to the standard notion $\ell_{\textrm{id}}$-online learnability, and we see that for any $x\in[0,1]$,
@@
so the result follows by Lemma 9. ∎
@@
Corollary 6.
If for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable in the realizable case, or for every $\epsilon>0$, ${\mathcal{H}}$ is $L_{\epsilon}$-online learnable in the realizable case, then the same holds of the randomization class of ${\mathcal{H}}$, and thus for both the distribution class and the dual distribution class.
@@
Fact 28 (Attias et al. (2023, Theorem 4)).
@@
Proof.
We first show that $\mathrm{Onl}_{\ell}({\mathcal{H}})\geq d\cdot\ell(\gamma)$. This is only a slight modification of Lemma 8.
@@
where $t_{i}$ is the restriction of $b$ to level $i$. As $\tau$ takes a constant value $\ell(\gamma)-\epsilon$, the online dimension is greater than $d(\ell(\gamma)-\epsilon)$. ∎
@@
Proof.
Here we will use the connection of agnostic online learnability with stability, which was utilized earlier in the appendix.
@@
For a base case, let $d=1$. We simply need that $X$ and ${\mathcal{H}}^{\prime}$ are nonempty, which is satisfied by the existence of any shattered tree.
Now assume that $d$ is such that we have the inductive invariant for $d$: if ${\mathcal{H}}^{\prime}$ is a class on $X$ that $\epsilon$-shatters a binary tree of depth $\frac{k^{d+1}-1}{k-1}$, then there are $x_{1},\dots,x_{d}\in X$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}^{\prime}$ such that for all $i<j$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$.
Also assume the hypothesis of the invariant for $d+1$: ${\mathcal{H}}^{\prime}$ is a class on $X$ that $\epsilon$-shatters a binary tree $T$ of depth $\frac{k^{d+2}-1}{k-1}$. As before, by the width of a real interval with endpoints $a<b$, we mean $b-a$. We pick some $h\in{\mathcal{H}}^{\prime}$, partition $[0,1]$ into $k$ intervals $I_{1},\dots,I_{k}$ each of width at most $\epsilon-2\delta$, and then partition $X$ into sets $X_{1},\dots,X_{k}$ where if $x\in X_{i}$, then $h(x)\in I_{i}$. Then by our Ramsey result for trees, Corollary 4, as $k\left(\frac{k^{d+1}-1}{k-1}+1\right)-k+1=\frac{k^{d+2}-1}{k-1}$, the depth of $T$, some $X_{a}$ contains the set of values decorating a subtree $T^{\prime}$ of $T$ of depth $\frac{k^{d+1}-1}{k-1}+1$. Let $x^{\prime}$ be the root of $T^{\prime}$.
Let ${\mathcal{H}}_{L}\subseteq{\mathcal{H}}^{\prime}$ consist of all hypotheses $\Lambda(b)$ where $b$ is a branch of $T$ extending $x^{\prime}$ to the left, and let ${\mathcal{H}}_{R}\subseteq{\mathcal{H}}^{\prime}$ consist of all hypotheses $\Lambda(b)$ where $b$ is a branch of $T$ extending $x^{\prime}$ to the right. By the spread-shattering hypothesis, if $h_{L}\in{\mathcal{H}}_{L}$ and $h_{R}\in{\mathcal{H}}_{R}$, then $|h_{L}(x^{\prime})-h_{R}(x^{\prime})|\geq\epsilon$. Because $I_{a}$ has width at most $\epsilon-2\delta$, either the set $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$ or the set $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$ has distance to $I_{a}$ at least $\delta$. Assume without loss of generality that it is $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$. The tree of left descendants of $x^{\prime}$ in $T^{\prime}$ has depth $\frac{k^{d+1}-1}{k-1}$, and is $\epsilon$-shattered by ${\mathcal{H}}_{L}$. Thus by the inductive hypothesis, there are left descendants $x_{1},\dots,x_{d}$ of $x^{\prime}$ in $T^{\prime}$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}_{L}$ for each $i$ such that for each $i<j\leq d$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$. We now let $x_{d+1}=x^{\prime}$ and $h_{d+1}=h$, and observe that for $i\leq d$, $h_{i}\in{\mathcal{H}}$ while $h(x_{i})\in I_{a}$, so $|h_{i}(x^{\prime})-h(x_{i})|\geq\delta$. ∎
@@
Proof.
Many of these implications follow from Lemma 10. To show all of the agnostic case conditions are equivalent, it suffices to show that if for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable, ${\mathcal{H}}$ is online learnable.
@@
Using Lemma 10 once more, it suffices to show that if for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable in the realizable case, ${\mathcal{H}}$ is online learnable in the agnostic case. The contrapositive of this follows from the other part of Lemma 11. ∎
7 Complexity of prediction for statistical classes
Thus far we considered the sample complexity needed to learn. To learn a real-valued function we need to perform empirical risk minimization: find the $h\in{\mathcal{H}}$ that minimizes the empirical risk, or is within a given $\epsilon$ of the minimum.
In our case, the function $h$ we are learning usually has an infinite domain: it maps parameters to the corresponding measures. Thus it is not clear how to represent it. Further, there can be many $h$’s that minimize empirical risk on a sample set, and they may provide different predictions on other inputs. At test time, what we actually need is not $h$ itself, but the predictions generated by $h$. Thus, rather than pick some representation of an empirical risk minimizer and study how to generate it, in this section we will look at decision problems associated with the possible predictions of an empirical risk minimizer. We formalize this below.
Fix a real-valued function class ${\mathcal{H}}$ over a range space whose elements can be effectively represented – like the integers or rationals. Given a finite collection of pairs $(x_{1},y_{1})\ldots(x_{n},y_{n})$, a hypothesis in ${\mathcal{H}}$ is $\epsilon$-optimal if its error, using the standard mean squared loss, is within $\epsilon$ of the infimum of the error for any hypothesis in the class. It is $\epsilon$-fitting if its error is at most $\epsilon$. We can also drop $\epsilon$ and talk about optimal or fitting hypotheses for a training set, noting that such a hypothesis might not exist. We can consider prediction over $\epsilon$-optimal hypotheses:
Given a sequence of input-output pairs $x_{1},y_{1}\ldots x_{n},y_{n}$, where $y_{i}$ are rational, along with a new input $x$, a rational $q$, and rational $\epsilon\geq 0$, determine whether for each $h\in{\mathcal{H}}$ that is $\epsilon$-optimal, $h(x)\geq q$.
We can likewise consider $\epsilon$-fitting, optimal, fitting. Notice that these definitions do not require us to represent the functions: instead we quantify over their predictions. We will be interested in the complexity of this problem for a fixed ${\mathcal{H}}$, $\epsilon$, and $q$: thus the input is only $x_{1},y_{1}\ldots x_{n},y_{n}$.
When ${\mathcal{H}}$ is the randomization class of a concept class or function class, it is not obvious how to represent elements of the range space – the inputs $x_{i}$ – since they are random variables. We focus on the case of the dual distribution function class formed over a hypothesis class ${\mathcal{H}}_{0}$. Recall that the parameter space of this class consists of measures on the range space of ${\mathcal{H}}_{0}$, and each such distribution $\mu$ specifies a function that mapping $h_{p}\in{\mathcal{H}}_{0}$ ($C_{p}$ for a concept class) to its $\mu$ mean (resp. its $\mu$ probability). For a concept class, our training set consists of elements of the parameter space and the corresponding $\mu$-probability, while in the real-valued case we have the corresponding $\mu$-mean.
We look first at the case where ${\mathcal{H}}_{0}$ comes from a definable family over a first-order structure. This will allow us to relate the decidability of the prediction problem to decidability of satisfaction for logics over the structure. We always assume that the probabilities in our training set, as well as the tolerance parameter $\epsilon$, are presented as rationals, coded in binary. We likewise restrict training inputs – which are elements of the model – to be given by constants of the model: in the case of models over the reals, they would be rational.
It is easy to show that for decidable structures – those where we can decide whether a first-order sentence is true – any of these prediction problems are decidable, not just for the definable families themselves, but for the distribution class families. For some common decidable structures, we can even get reasonable bounds in terms of the size of the training set.
{restatable}
theoremptimeerm If $\mathfrak{M}$ is a decidable first-order structure, then the optimal prediction problems for the dual distribution or distribution class are decidable as well. Let $\mathfrak{M}$ be any of Presburger arithmetic, the real closed field, the real closed order group. The $\epsilon$-fitting prediction problem can be solved in NP for the distribution class (or the dual distribution class) formed over any fixed ${\mathcal{C}}_{\phi}$, and in PTIME when fixing $\epsilon=0$.
Note that since we are fixing the partitioned formula $\phi(\vec{x};\vec{y})$, we are fixing the number of parameters that need to be learned. This contrasts with most prior work on fitting problems (Abrahamsen et al., 2021; Bertschinger et al., 2023; Goel et al., 2021), which provide hardness results in settings where the number of weights is part of the input.
Before turning to statistical classes, we show that the optimal prediction problem is decidable as long as the underlying structure is decidable:
Proposition 8.
If ${\mathcal{H}}_{0}$ is a concept class given by partitioned formula $\phi(\vec{x};\vec{y})$ in a decidable model, then all the prediction problems (fitting, optimal, $\epsilon$-optimal) are decidable for ${\mathcal{H}}_{0}$.
Here we assume decidability of arbitrary sentences for simplicity, but it will be clear from the proof that we only need to decide certain formulas related to $\phi$.
Proof.
We assume the realizable case – the optimal hypothesis gives error zero – for brevity.
If the defining formula is $\phi(\vec{x},\vec{y})$, our sequence of inputs will be $\vec{x}_{1}\ldots\vec{x}_{n}$, and our outputs $y_{i}$ will be 00 or $1$. It suffices to show that the problem is decidable when the loss is 00 – since we can enumerate all subsets $S$ of $1\ldots n$ and look for hypotheses that are correct on this subset. For simplicity assume that our sample data asserts that $h(\vec{x}_{i})$ is true on each $i$. Since the output of $h$ on the new element $x_{n+1}$ is binary, our loss on $x_{n_{1}}$ is discrete, and we can consider the two possible outputs separately. To determine if the outcome true holds on input $x_{n+1}$, we need to decide if:
$\forall\vec{y}~{}\left[\bigwedge_{i\in S}~{}\phi(\vec{x}_{i},\vec{y})\rightarrow\phi(\vec{x}_{n+1},\vec{y})\right]$
We can do this appealing to the decidability of the structure. ∎
Of course, we are not interested in prediction problems about the original hypothesis class ${\mathcal{H}}_{\phi}$. Instead we are interested in prediction problems for statistical classes based on ${\mathcal{H}}_{\phi}$. We next turn to the dual distribution class, and show decidability of the optimal prediction problem:
Proposition 9.
If ${\mathcal{H}}_{\phi}$ is the dual distribution class built over a concept class given by a partitioned first-order formula $\phi(\vec{x};\vec{y})$ in a decidable model, then the optimal prediction problems are decidable.
Here we deal with the dual distribution class – where the measures are defined over the range space $\vec{x}$. But the same comment holds for the distribution class, just swapping the roles of $\vec{x}$ and $\vec{y}$.
Proof.
For simplicity, we only give the argument for the $\epsilon$-fitting version. In this case our decision problem reduces to:
Given $\epsilon$, $\vec{y}_{i},r_{i}:i\leq n$, $\vec{y}_{n+1}$ and $q$, is there a measure $\mu$ such that $\Sigma_{i\leq n}|\mu(\vec{x}|\phi(\vec{x},\vec{y}_{i}))-r_{i}|<\epsilon$ and $\mu(\vec{x}|\phi(\vec{x},\vec{y}_{n+1}))>q$.
All that is relevant from $\mu$ is its values on the atomic measure algebra generated by $\phi(\vec{x},\vec{y}_{i})$ for $i\leq n+1$. Let $A$ be the generators of this algebra, which have size at most $2^{n+1}$. Our algorithm will have two stages, the decision procedure stage, and the linear arithmetic stage. In the decision procedure stage, we make $2^{n+1}$ calls to the decision procedure for the theory to determine the subset $A_{\emptyset}$ of Boolean combinations that are empty. Let $A_{i}$ be the elements of the algebra that include a conjunct $\phi(\vec{x},\vec{y}_{i})$. We are thus asking whether there are numbers $r_{a}\in[0,1]$ for $a\in A$ such that:
  • $\Sigma_{a\in A\setminus A_{\emptyset}}r_{a}=1$
  • For $i\leq n$, $\Sigma_{a\in A_{i}\setminus A_{\emptyset}}|r_{a}-r_{i}|<\epsilon$
  • $\Sigma_{a\in A_{n+1}\setminus A_{\emptyset}}r_{a}>q$.
In the linear arithmetic stage we can solve this by appealing to the decidability of the reals with addition and order (Weispfenning, 1988). In the above case, it suffices to invoke decidability of solving systems of linear inequations. ∎
Above we did not make any assumption on the base hypothesis class in terms of learnability/VC dimension/fat-shattering. Note that if the model is NIP, then the size of $A-A_{\emptyset}$ will be polynomial in $n$, say $O(n^{k})$. We can compute $A-A_{\emptyset}$ with $O(n^{k})$ many calls to the decision procedure. We iteratively find elements of $A-A_{\emptyset}$ represented by Boolean combinations of at most $i$ elements. At round $i+1$ we take all the combinations of size $i$ and make calls to the decision procedure to see if the intersection with $\phi(\vec{x},\vec{y}_{i+1})$ and its complement are both non-empty, adding any non-empty element to the working set. Our working set never grows past $O(n^{k})$, and thus at each of the $n$ rounds we make at most $O(n^{k})$ calls.
Thus the decision procedure stage consists of polynomial many calls to a decision procedure for the model. However, each of these calls concerns a formula of the form $\Gamma(U_{1},U_{2})=\exists\vec{x}~{}\bigwedge_{\vec{a}_{i}\in U_{1}}\phi(\vec{x},\vec{a}_{i})\wedge\bigwedge_{\vec{a}_{i}\in U_{2}}\neg\phi(\vec{x},\vec{a}_{i})$, where $U_{1},U_{2}$ will be finite sets whose size will not grow in the process.
The comments above apply to the dual distribution class or the distribution class equally: since one has finite VC-dimension iff the other does, and since $\phi$ is fixed in our complexity analysis any blow up in moving between a class and its dual is irrelevant.
To complete the proof of the complexity assertions in Theorem 7, we show that for common decidable models like those listed in the theorem, the linear programming phase can be handled in polynomial time in the sizes of $U_{1}$ and $U_{2}$, for fixed $\phi$. This follows easily from:
Theorem 11.
Let $M$ be any of Presburger arithmetic, the real closed field, the real closed order group. For any fixed $\Gamma$, there is an algorithm that decides statements of the form $\Gamma(U_{1},U_{2})$ in polynomial time in the size of $U_{1},U_{2}$.
Proof.
We rely on two properties of the model. One property is that it has quantifier elimination in a language where atomic formulas can be implement in polynomial time: this property is well-known for the models in the theorem. The second property is restricted quantifier collapse, which is also known for these models: we explain this property next.
Let $U$ be a finite relational vocabulary – for the application, $U=\{U_{1},U_{2}\}$ suffices. A restricted quantifier sentence is a sentence in the language of $L(M)\cup U$ built up from atomic formulas via Boolean operators and quantifications over predicates in $U$. In this case $\exists x\in U_{i}~{}\phi$. It is known (see e.g. (Benedikt, 2006)) that in each of the models above, we can convert an arbitrary $L(M)\cup U$ sentence into a restricted-quantifier sentence. Clearly each such sentence can be evaluated in time polynomial in the sizes of predicates in $U$. ∎
Thus the only source of intractability is in the linear programming part. In the case $\epsilon=0$ we are simply determining feasibility of a set of linear equations. Thus using standard algorithms for feasibility of linear systems (Schrijver, 1986) (e.g. ellipsoid), we conclude that the problem is in polynomial time. In the general case it is in NP.
Extensions to real-valued classes. We now discuss optimal prediction problems for distribution classes where the base class consists of real-valued functions.
Proposition 10.
Let ${\mathcal{H}}_{\phi}$ be a real-valued class given by a bounded family of real-valued functions definable by partitioned formula $\phi(\vec{x},y;\vec{p})$ over the real ordered field.
Then the $\epsilon$-optimal prediction problem for the distribution class or dual distribution class is decidable.
Proof.
We deal with the dual distribution class for simplicity. Consider parameters $\vec{p}_{1}\ldots\vec{p}_{n}$, let $\vec{F}=F_{1}\ldots F_{n}$ be the corresponding functions where $F_{i}$ is formed by fixing $\vec{p}_{i}$ in $\phi$. For a probability distribution $\mu$ such that the $\vec{F}$ are measurable, let $\textbf{Mean}_{\mu}(\vec{F})$ be the vector of $\mu$-means of $\vec{F}$: a vector in ${\mathcal{R}}^{n}$. Let $\textbf{Means}(\vec{F})$ be the collection of such vectors.
We claim that $\textbf{Means}(\vec{F})$ is definable over the real ordered field. For a fixed $j$, let $\textbf{Means}_{j}(\vec{F})$ be all the vectors that arise from measures that are convex combinations of $j$ point masses. Clearly this set is definable, since these are restricted $j$-sums of combinations of the ranges of the $F_{i}$. However, $\textbf{Means}(\vec{F})$ is the same as $\textbf{Means}_{n}(\vec{F})$: for discretely-supported distributions, this follows from Caratheodory’s theorem (Cook & Webster, 1972). While for general distributions we will show belwow that it can reduced to the finitely-supported case. Note that this argument shows that in the concept class case, without assuming NIP, we can always restrict to probability distributions that give non-zero mass to only polynomially many Boolean combinations of the concepts in the training set. However, in the NIP case, we can compute a superset of these combinations without considering the training output values.
All of our prediction problems reduce to statements about $\textbf{Means}(\vec{F})$, and since $M$ is decidable, we can decide these.
We now fill in the details about why $\textbf{Means}(\vec{F})$ is the same as $\textbf{Means}_{n}(\vec{F})$ for general distributions. First consider measures $\mu$ is an arbitrary measure on some arbitrary domain $X$. Let $F_{1}\ldots F_{k}$ be arbitrary bounded real-valued valued functions on $X$.
First, note that by the law of large numbers, for every number $n$, there is a finitely supported probability measure $\mu^{n}$ such that for each $i\leq k$, the $\mu$-mean of each $F_{i}$ is within $\frac{1}{n}$ of the $\mu^{n}$ mean of $F_{i}$. By Caratheodory’s theorem, we can take each $\mu^{n}$ to have support of size $k$, thus given by $k$-tuple of domain elements $\vec{x}^{n}$ and a $k$-tuple of reals $\vec{r}^{n}$ in $(0,1]$, with $\vec{r}^{n}$ summing to one. By moving to a subsequence, we can assume $\vec{r}^{n}$ converges in the product topology over the reals. Let $r_{j}$ be the limit of the $j^{th}$ component.
Now suppose that the domain $X$ of our functions, and of our measure $\mu$ is a product of the reals. For $j\leq k$ let $\vec{x}^{n}_{j}$ denote the $j^{th}$ component of $\vec{x}^{n}$.
Let $D$ be the set of $j$ such that $\vec{x}^{n}_{j}$ does not eventually stay in some compact subset of the reals: thus $D$ is the set of “divergent indices”. Since $\mu$ is a probability distribution, the mass of $\mu$ outside of compact set must go to zero as the compacts expand towards infinity. From this we see that for $j\in D$, the weight $r^{n}_{j}$ corresponding to $x^{n}_{j}$ must go to zero. Thus $r_{j}=0$ for $j\in D$, which means that the sum of $r_{j}$ for $j$ outside of $D$ must be $1$.
For $j$ outside of $D$, $\vec{x}^{n}_{j}$ is eventually contained in a compact set, so by moving to a subsequence we can assume it converges to some real vector $\vec{x}_{j}$.
Up until this point, we have not used any additional properties of the $F_{i}$. Let us now assume that $F_{i}$ are continuous. Then for each $j$ outside of $D$ $F_{i}(\vec{x}^{n}_{j})$ converges to $F_{i}(\vec{x}_{j})$.
It is now clear that, again assuming continuity of each $F_{i}$, if we take the measure $\mu^{\infty}$ that assigns $\vec{x}_{j}$ to $r_{j}$ for $j\leq k$ outside of $D$, we have the finitely-supported probability measure we need.
We note that this argument for continuous $F_{i}$ does not require the measure to be a probability distribution, but applies also to a subprobability distribution.
Finally, we will make use of the fact that our functions $F_{i}$ are not arbitrary function. Recall that in the theorem we have a formula $\phi(\vec{x},y;\vec{p})$ first-order over the real field, such that for each $\vec{p}_{i}$, $\phi(\vec{x},y,\vec{p}_{i})$ defines a function, which we can denote by $F^{\phi}_{\vec{p}_{i}}$. Our functions $F_{i}$ are all $F^{\phi}_{\vec{p}_{i}}$ for some $\vec{p}i$.
We utilize the following fact about definable functions, which holds for every $o$-minimal expansion of the real field:
Fact 29.
van den Dries (1998) For every $\phi(x_{1}\ldots x_{d},y;\vec{p})$ there is a number $j$ such that for each $\vec{p}_{0}$, there is a partition of ${\mathcal{R}}^{d}$ into definable sets $S_{1}\ldots S_{j}$ such that $F^{\phi}_{\vec{p}_{0}}$ is continuous on each $S_{i}$.
By taking a common refinement, we can assume a single partition for each of our functions $F^{\phi}_{\vec{p}_{i}}$. Thus, by the argument for continuous functions above, we can get finitely-supported measures that correctly produce the mean of each function on each partition. By taking the weighted sum of these measures, weighting the measure for $S_{i}$ by the $\mu$ measure of $S_{i}$ we can get a single finitely-supported measure.
Note that the process above gives a measure with support larger than the number of functions. But applying Caratheodory’s theorem again, we can reduce the support to equal the number of functions. ∎

### v2

@@ 5.4 An Alternate Approach to Realizable Online Learning
Thus far we have shown that realizable online learning is not preserved under moving to statistical classes. We have indicated that this is related to another pathology, that online learnability is not preserved under applying continuous mappings. We will now look at using this intuition to “fix” the pathologies of realizable online learning. We will do this by changing the definition, applying alternate loss functions which do not sum the cumulative losses, but rather discretize them.
@@
Definition 53 (Online learnability for a general loss function)
@@
Proof  The same learning algorithm will suffice. On any given run of the algorithm, the regret with $\ell_{1}$ as the loss function will be at most the regret with $\ell_{2}$ as the loss function:
@@
so regret with loss function $\ell_{1}$ will satisfy the same upper bound required of regret with loss function $\ell_{2}$, up to the fixed factor $C$.
@@
Definition 55 ($\epsilon$-truncated loss functions)
@@
Proof  Online learnability is equivalent to the standard notion $\ell_{\textrm{id}}$-online learnability, and we see that for any $x\in[0,1]$,
@@
so the result follows by Lemma 54.
@@
Corollary 58
If for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable in the realizable case, or for every $\epsilon>0$, ${\mathcal{H}}$ is $L_{\epsilon}$-online learnable in the realizable case, then the same holds for both the distribution class and the dual distribution class.
@@
Fact 26 (Attias et al. (2023, Theorem 4))
@@
Proof  We first show that $\mathrm{Onl}_{\ell}({\mathcal{H}})\geq d\cdot\ell(\gamma)$. This is only a slight modification of Lemma 49.
@@
where $t_{i}$ is the restriction of $b$ to level $i$. As $\tau$ takes a constant value $\ell(\gamma)-\epsilon$, the online dimension is greater than $d(\ell(\gamma)-\epsilon)$.
@@
Proof  Here we will use the connection of agnostic online learnability with stability, which was utilized earlier in Section 5.2.
@@
Consider the base case, $d=1$. We simply need that $X$ and ${\mathcal{H}}^{\prime}$ are nonempty, which is satisfied by the existence of any shattered tree.
Now assume that $d$ is such that we have the inductive invariant for $d$: if ${\mathcal{H}}^{\prime}$ is a class on $X$ that $\epsilon$ spreads-shatters a binary tree of depth $\frac{k^{d+1}-1}{k-1}$, then there are $x_{1},\dots,x_{d}\in X$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}^{\prime}$ such that for all $i<j$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$.
Also assume the hypothesis of the invariant for $d+1$: ${\mathcal{H}}^{\prime}$ is a class on $X$ that $\epsilon$ spread-shatters a binary tree $T$ of depth $\frac{k^{d+2}-1}{k-1}$. As before, by the width of a real interval with endpoints $a<b$, we mean $b-a$. We pick some $h\in{\mathcal{H}}^{\prime}$, partition $[0,1]$ into $k$ intervals $I_{1},\dots,I_{k}$ each of width at most $\epsilon-2\delta$, and then partition $X$ into sets $X_{1},\dots,X_{k}$ where if $x\in X_{i}$, then $h(x)\in I_{i}$. Then by our Ramsey result for trees, Corollary 47, as $k\left(\frac{k^{d+1}-1}{k-1}+1\right)-k+1=\frac{k^{d+2}-1}{k-1}$, the depth of $T$, some $X_{a}$ contains the set of values decorating a subtree $T^{\prime}$ of $T$ of depth $\frac{k^{d+1}-1}{k-1}+1$. Let $x^{\prime}$ be the root of $T^{\prime}$.
Let ${\mathcal{H}}_{L}\subseteq{\mathcal{H}}^{\prime}$ consist of all hypotheses $\Lambda(b)$ where $b$ is a branch of $T$ extending $x^{\prime}$ to the left, and let ${\mathcal{H}}_{R}\subseteq{\mathcal{H}}^{\prime}$ consist of all hypotheses $\Lambda(b)$ where $b$ is a branch of $T$ extending $x^{\prime}$ to the right. By the spread-shattering hypothesis, if $h_{L}\in{\mathcal{H}}_{L}$ and $h_{R}\in{\mathcal{H}}_{R}$, then $|h_{L}(x^{\prime})-h_{R}(x^{\prime})|\geq\epsilon$. Because $I_{a}$ has width at most $\epsilon-2\delta$, either the set $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$ or the set $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$ has distance to $I_{a}$ at least $\delta$. Assume without loss of generality that it is $\{h_{L}(x^{\prime}):h_{L}\in{\mathcal{H}}_{L}\}$. The tree of left descendants of $x^{\prime}$ in $T^{\prime}$ has depth $\frac{k^{d+1}-1}{k-1}$, and is $\epsilon$-shattered by ${\mathcal{H}}_{L}$. Thus by the inductive hypothesis, there are left descendants $x_{1},\dots,x_{d}$ of $x^{\prime}$ in $T^{\prime}$ and $h_{1},\dots,h_{d}\in{\mathcal{H}}_{L}$ for each $i$ such that for each $i<j\leq d$, $|h_{i}(x_{j})-h_{j}(x_{i})|\geq\delta$. We now let $x_{d+1}=x^{\prime}$ and $h_{d+1}=h$, and observe that for $i\leq d$, $h_{i}\in{\mathcal{H}}$ while $h(x_{i})\in I_{a}$, so $|h_{i}(x^{\prime})-h(x_{i})|\geq\delta$.
@@
Proof  Many of these implications follow from Lemma 56. To show all of the agnostic case conditions are equivalent, it suffices to show that if for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable, ${\mathcal{H}}$ is online learnable.
@@
Using Lemma 56 once more, it suffices to show that if for every $\epsilon>0$, ${\mathcal{H}}$ is $\ell_{\epsilon}$-online learnable in the realizable case, ${\mathcal{H}}$ is online learnable in the agnostic case. The contrapositive of this follows from the other part of Lemma 60.